if pam can paint a car in 4 hours and pat can paint the same in 6 hours how long will it take for both to paint the car together?
Answer by jorel1380(3719)

t/4+t/6=1
3t+2t=12
5t=12
t=12/5
It takes 12/5ths of an hour to paint the house together..